OpenStudy (anonymous):

10 gal. bucket of product will make 100 lbs. , which will fill 1 cubic yard. How much material is needed to fill 600 ft^2 area 2 inches thick?

jigglypuff314 (jigglypuff314):

so 10gal = 1yd^3 = 9ft^3 you are given 600ft^2 * 2in 2in = 1/6 ft so 600ft^2 * 1/6 ft = 100ft^3 then to get to gal. multiply 100ft^3 by (10/9)